OpenStudy (anonymous):

How do I write: x^2-6x+10=3 in (ax+b)^2=c form?

OpenStudy (anonymous):

get familiar with the fact that:\[(ax + b)^2 = a^2x^2 + 2axb + b^2\] now find the corresponding values of a and b. what's left over is c.

OpenStudy (anonymous):

Okay. So in this case, from the equation, which values are what in the formula?

OpenStudy (anonymous):

a = 1 b = -3

OpenStudy (anonymous):

What? I didn't mean the answer, I meant from: \[x ^{2}-6x+10=3\]Which values from this make up the formula?